OpenStudy (anonymous):

Solve for x in the proportion:

OpenStudy (anonymous):

\[\frac{ x + 1 }{ 8 }\] = \[\frac{ 2x - 5 }{ 2 }\]

OpenStudy (anonymous):

I got 2x + 1 = 16x - 40 Not sure what to do next

OpenStudy (johnweldon1993):

You mean you got 2x + 2 = 16x - 40 right?

OpenStudy (anonymous):

Yes, sorry typo.

OpenStudy (anonymous):

Do you think the answer is 3? It's a choice. and I plugged in. 8 = 8

OpenStudy (johnweldon1993):

No problem....well now...you want to combine like terms...you want every number with an 'x' on one side of the equation...and every other number on the other side....so subtract 2x from both sides and add 40 to both sides 2x + 2 = 16x - 40 -2x -2x 2 = 14x - 40 +40 +40 42 = 14x divide both sides by 14 to solve for x 42 = 14x --------- 14 14 x = 3

OpenStudy (anonymous):

Ahh I see, well these problems I struggle with. Thank you for your explanation (:

OpenStudy (johnweldon1993):

No problem! :)